Find x and y: x^y=y^x and x=2y

Find x and y:
x^y=y^x   and   x=2y

Grade:11

1 Answers

Rinkoo Gupta
9 years ago
x=2y=> y=x/2 put this value in eq(1)
xy=yx
x(x/2)=(x/2)x
=> x=[(x/2)x]2/x
=> x=(x/2)2
=> x=x2/4
=>4x=x2
=>4x-x2=0
=>x(4-x)=0
=> x=0,4
Thanks & Regards
Rinkoo Gupta
AskIITians Faculty